How do I solve by elimination: X-4Y=-23
2X+3Y=9

Answer by stanbon(75887)

How do I solve by elimination:
1st: X-4Y=-23
2nd: 2X+3Y=9
---------
Multiply 1st by 2 to get:
3rd: 2X-8Y=-46
---------
Subtract 2nd from 3rd to solve for "Y":
-11Y = -55
Y = 5
---------
Substitute into X-4Y=-23 to solve for X:
X-4*5 = -23
X = -3
------------
Checking answer in 2X+3Y=9
2*-3+3*5 = 9
-6+15=9
9=9
========
Cheers,
Stan H.